# Depth-constrained ASV navigation with deep RL and limited sensing

### v2

@@ III-C GP depth estimation
$d(\mathbf{x})\sim\mathcal{GP}(\mu_{0}(\mathbf{x}),k(\mathbf{x},\mathbf{x}^{\prime})),\$ (3)
@@ III-E POMDP Formulation
A Partially Observable Markov Decision Process (POMDP) is a tuple $(S,A,O,P,Z,R,\gamma)$, where $S$ is a set of partially observable states; $A$ is a set of actions; $Z$ is a set of observations; $P$$S\times A\rightarrow\Pi(S)$ is the state-transition model, mapping to a probability distribution $\Pi(\cdot)$ over states; $O$$S\times A\rightarrow\Pi(Z)$ is the observation model; $R$ is the reward function and $\gamma\in[0,1)$ is a discount factor. The probability distribution over states $\pazocal{B}=\Pi(S)$, called belief, is used to model uncertainty about the true state. The goal of solving a POMDP, e.g., with RL, is to compute a policy $\pi:\pazocal{B}\rightarrow A$, to maximize the expected return:
@@ III-E4 Reward function
$r_{\text{depth}}=\begin{cases}0&\text{if }\frac{1}{3}L_{\text{d}}\leq z_{t}<\frac{2}{3}L_{\text{d}}\\ \frac{3}{L_{d}}\left(z_{t}-\frac{2}{3}L_{\text{d}}\right)&\text{if }\frac{2}{3}L_{\text{d}}\leq z_{t}<L_{\text{d}}\\ \frac{3}{L_{d}}\left(z_{t}-\frac{1}{3}L_{\text{d}}\right)&\text{if }0\leq z_{t}<\frac{1}{3}L_{\text{d}}\\ \end{cases}$ (13)
where $z_{t}$ is the depth measurement at time $t$ and $L_{d}$ is the depth limit that defines the unsafe regions. In this formulation, no penalty is applied when the depth lies within the safe range, $[\frac{1}{3}L_{d},\frac{2}{3}L_{d})$. However, as the depth measurement approaches either the lower or upper bounds (i.e., deviates from the safe interval), the penalty increases linearly. An illustration of this reward function is presented at Figure 5. This design effectively encourages the agent to maintain a safe operating depth, thereby avoiding proximity to both the shore and excessively deep regions.
@@ IV-A Implementation and training details
[…]

### v3

@@ III-C GP depth estimation
$d(\mathbf{x})\sim\mathcal{GP}(\mu_{0}(\mathbf{x}),k(\mathbf{x},\mathbf{x}^{\prime})),\\$ (3)
@@ III-E POMDP Formulation
[…]
@@ III-E4 Reward function
[…]
where $z_{t}$ is the depth measurement at time $t$ and $L_{d}$ is the depth limit that defines the unsafe regions. In this formulation, no penalty is applied when the depth lies within the safe range, $\frac{1}{3}L_{d},\frac{2}{3}L_{d})$. However, as the depth measurement approaches either the lower or upper bounds (i.e., deviates from the safe interval), the penalty increases linearly. An illustration of this reward function is presented at Figure 5. This design effectively encourages the agent to maintain a safe operating depth, thereby avoiding proximity to both the shore and excessively deep regions.
@@ IV-A Implementation and training details
We generate 120 depth maps in simulation for training (Figure 3), randomizing the shoreline and depth values with radial gradients, noise and filtering techniques. The maps are sized $50\times 50$ $\mathrm{m}$ with a depth resolution of $10\text{\,}\mathrm{c}\mathrm{m}$2. We used the dynamic model of the ASV and implement a vectorized OpenAI gym environment, which allows for simulating multiple ASV’s in parallel, achieving approximately 2000 FPS on a single GeForce 4070 GPU. The RL environment is implemented using the Stable-Baselines3 library [30], and the safe environment is constructed with the OmniSafe library [31]. For our proposed method, we employ a two-layer multilayer perceptron (MLP) with 256 units per layer. We train our agent using Soft Actor-Critic (SAC) [32], which is a popular off-policy reinforcement learning algorithm that maximizes reward while optimizing for entropy, encouraging exploration and robust policies. For safe-RL, we used the Lagrangian version of SAC (SACLag) [33]. For the kernel function, we used the well-known Radial Basis Function (RBF) kernel [34] with a length-scale set to 3.0 empirically. We trained all the models for 5 million steps and for 5 random seeds. We evaluated approaches using five metrics: Success Rate (SR) measuring task completion percentage, Efficiency Score (ES) calculated as success rate divided by episodic length, Depth Break (DB) indicating episodes terminated by depth constraint violations, Velocity Smoothness (VS) quantifying velocity change smoothness, and Heading Smoothness (HS) measuring heading direction change rates [35].